OpenStudy (anonymous):

Solve the system using substitution: x^2+y^2=32 y=x How do i do this?

OpenStudy (anonymous):

substitute x in for y then add together to get 2x^2 = 32 solve from there

OpenStudy (anonymous):

substitute y=x in the first equation x^2+x^2=32 2x^2=32 x^2=16 x=(plus or minus)4=y

OpenStudy (anonymous):

Multiple choice answers are: (-4,-4) (-4,4) (4,-4) (4,4) which one would it be?

OpenStudy (anonymous):

(-4,-4) and (4,4)

OpenStudy (anonymous):

Thank you so much. What about: x^2-y+5=0 x-y+5=0